# The DAMIC-M Low Background Chamber

### v1

@@ 2 Prototype Objectives
First results on dark matter-electron scattering [24], including a search for a daily modulation in the event rate [31], were published, and studies of alternative low-mass dark matter candidates are forthcoming. In addition, three doctoral theses describe work on the LBC [44, 45, 46].
@@ 3 Detector Description
The LBC is designed to provide $O(\mathrm{1})$ dru background environment for the operation of CCDs with modular electronics. The innermost region of the detector consists of a cold copper housing, which simultaneously serves to cool the CCDs and provide an infrared shield from the surrounding room temperature materials. This cold copper housing is enclosed by 2 cm of ancient Roman lead from a sunken galley [47, 38], then an additional 4-10 cm of very low background ("très faible activité" or TFA) lead. Kinked channels provide a pathway for the two-layer polyimide flex cables, used for CCD operation and readout, to exit the inner region eliminating a line-of-sight to the inner copper housing. Cold oxygen-free high-conductivity, or OFHC [^2] In reality, it is from OFE-OK copper grade by Luvata in Finland. Because it is comparable to oxygen-free high thermal conductivity copper used in other experiments, we will use acronym OFHC., copper "fingers" travel through four narrow gaps in the interior (warm) lead shielding with $\sim$1 mm clearance to provide thermal contact and mechanical support for the inner copper housing.
Outside of the interior lead shielding are four axially-spaced intermediate boards, which provide the first stage amplification to the CCD video signal. The boards connect to flex cables through commercial 50-pin connectors on each side. The output side of each board connects to a 45 cm long, five-layer CCD flex cable that exits below the lead shielding to commercial 50-pin D-Sub vacuum feedthroughs below.
@@ 3.1 Charge-Coupled Devices
These 6k×$\times$4k CCD have two skipper amplifiers on one side and two standard amplifiers on the opposite side. Unlike conventional CCD amplifiers, skipper amplifiers can be configured to make multiple, non-destructive charge measurements (NDCMs) [25, 26, 27, 28]. Skipper readout essentially moves the charge contained in each pixel back-and-forth into the readout node allowing for many measurements of the same pixel so that they can be averaged. Since the measurements are uncorrelated, the readout noise is reduced to $\sigma_{1}/\sqrt{N_{\rm{skip}}}$, where $\sigma_{1}$ is the single-sample readout noise (the standard deviation of a single charge measurement) and $N_{\rm{skip}}$ is the number of NDCMs. By taking a large enough number of NDCMs, the readout noise can reach the sub-electron level and single charges [^4] For the sake of clarity, we use the term electrons to indicate charge carriers detected in the CCD. However, holes are held in the pixels of p-channel LBNL CCDs. are resolved. The single-electron resolution also provides a straightforward way of calibrating the energy response of the detectors.
A two-layer CCD flex cable carries the clock voltages to and video outputs from each CCD. The cable is glued to an extended non-functional area of the CCD. To improve adhesion, rectangular silicon shims were glued between the CCD die and the flex. The glue is low-background Epotek 301-2 previously selected for DAMIC at SNOLAB. The cable lines are wedge bonded with 25-$\mu$m thick Al wire to corresponding pads on the CCD. Special ESD safety precautions have been implemented during handling, packaging and testing CCDs to reach high packaging yield. During this time, care is taken to minimize the exposure of the CCDs to the air and hence, the accumulation of radon progeny on the CCD surface.
@@ 3.2 Boards and Cabling
A second-stage 45 cm long flexible cable, still thermally isolated from warm lead, goes from the intermediate board down to the cross part of the cryostat, where it is plugged into a two-to-one board. This board merges two cables coming from opposite sides of the same CCD and connects to the vacuum feedthrough. The two-to-one boards were needed only for the 6k×$\times$4k CCDs and were taken out when the CCD modules were installed in early 2023. The CCD modules have only one cable to carry all signals so new intermediate boards were also installed.
@@ 3.3 Cold Copper Parts and IR Shielding
The CCDs are housed in a copper box which cools the devices and acts as an IR shield. The dominant thermal path to the CCDs is through the CCD flex cables which are clamped to the CCD box via t-bars. In this configuration, the temperature remains stable when the on-chip CCD amplifiers are powered, see Sec. 6.1.
@@
The connection to the cold tip of a cryocooler is made by a flexible thermal strap by Thermal Management Technologies. Individual copper pieces were tightly connected with brass screws and a thin layer of thermally conductive grease (i.e. Apiezon N cryogenic high vacuum grease) was applied between them. For an electrical isolation, two 25 $\mu$m-thick Kapton tapes are placed between the cold tip and flexible strap.
@@
Because the copper lids are directly above and below CCDs, they are a major source of radioactive background. OFHC copper lids were used while taking science data in 2022 [24, 31], but were taken out later the same year. The new installed lids are from less radioactive, electro-formed copper (EFC) grown and machined by Laboratorio Subterraneo de Canfranc (LSC) [54]. The EFC lids not only have much lower radioactive contamination than the commercial OFHC copper, but also their cosmogenic activation time has been reduced from many months to less than a week, see Sec. 5.1.
@@ 3.5 External Passive Shielding and Support Structure
External gamma radiation and neutrons are attenuated by a layer of low background lead (faible activitè or FA) from LEMER and HDPE by Abaqueplast, both in France. The shielding around and above the cryostat cylinder, where the shielding of CCDs is the most important, consists of 15 and 20 cm of lead and HDPE, respectively, see Fig. 1 and 5. The bottom cross of the cryostat is surrounded by 5 cm of lead from all sides but below the cryostat and 10 cm of HDPE. The 41 cm diameter ring of lead shielding above the cryostat together with HDPE half-disks is supported by 2 cm thick stainless steel plate.
The support structure standing on the floor of the cleanroom is static, but motorized opening system can move half of the shield on rails to allow easy access to the cryostat. Neither the motor nor rails use lubrication and thus do not compromise the cleanroom cleanliness. A photo of fully assembled and partly opened detector is in Fig. 5.
@@ 3.6 LSM Infrastructure
The furniture and equipment inside the cleanroom are electrostatic discharge (ESD) safe to comply with requirements on handling ESD-sensitive equipment, like CCDs and electronics. While working with CCDs, we require the humidity above 35% and we deploy a local humidifier using deionized water if needed.
The radon concentration has been continuously measured every 10 minutes with Durridge RAD7 in the cleanroom since January 2023. The median radon concentration is (18.4$\pm$8.5) Bq/m3 with 99% of the data below 42 Bq/m3. Other backgrounds obtained by different groups are: gamma flux of 3.8$\times$10-6 cm-2s-1 between 4 and 6 MeV [56], thermal neutron flux of $4.6\times 10^{-6}$ n/cm2/s at the LBC location [57], and fast neutron flux above 1 MeV of (1.06$\pm$0.60)$\times$10-6 n/cm2/s [58].
@@ 4 Detector Readout and Instrumentation
The LBC is remotely controlled and monitored by custom data acquisition systems placed outside the external shield. We describe two different electronics systems installed in the LBC to run CCDs, of which the first one collected the data published in two DAMIC-M publications [24, 31]. The slow control, DAQ software and data quality monitor system are also detailed.
@@ 4.1 Readout Electronics
A custom data acquisition system based on a commercial Gen-III CCD controller from Astronomical Research Cameras, Inc. (ARC) was used for initial tests with the LBC. The controller provides the voltage biases and clocks required for operation of the CCD and reads out the video signal from two CCD amplifiers. External low-noise power supplies provide DC voltages for controllers, amplifiers and Vsub. Two CCD controllers were used to read two 6k×$\times$4k CCDs or two CCDs in two CCD modules.
@@
The new DAMIC-M readout electronics were successfully installed in the LBC in January 2024 and are used to control CCDs since then. These electronics were developed for DAMIC-M, have lower readout noise than the ARC CCD controller and can read all four CCDs in a module. The electronics chain consists of a custom-made Acquisition and Control Module (ACM) plugged to a VME crate (Wiener VME 6023x610), a 3 m long cable by Samtec and a custom-made front-end board. In addition, an adapter board connects the Samtec and D-Sub-50 connector in the vacuum feedthrough.
@@ 4.2 Instrumentation
The temperature of the cold copper is monitored by resistance temperature detectors (RTDs) PT-103 in two different positions. The cooling rate and temperature are set by a Lakeshore 336 controller that monitors the RTDs and controls the power to a 100 W heater installed next to the RTD closest to the cryocooler. To minimize temperature-induced stress on CCDs, we use the cooling rate of 0.1 K/min during temperature cycles. The system maintains the temperature within $\pm$0.2 K for many months, see Sec. 6.2.
@@
The detector and its components are protected by two-levels of backup power: (1) the DAMIC-M APC 2200 UPS and (2) the LSM system-wide UPS. In the case of a power outage, both UPS systems would be able to operate from battery power for at least one hour. This is adequate time for remote shifters to respond to alarms (see Sec 4.3) and safely shut down the experiment. In the unlikely scenario of an outage longer than one hour with no intervention from shift scientists, instruments would undergo a hard shut down.
@@ 4.3 Slow Control System
Conditions that can pose a hazard for the operation of the CCDs are triggered on by alarm daemons. For example, a failure of the vacuum pump would cause moisture to condense on the cold CCDs and introduce shorts in the electrical components. For this reason, the pump is powered by the local UPS which is monitored by the system. These alarms signal to on-shift users that the CCDs must be promptly shut down. In addition, a safety watchdog daemon is used to execute a shutdown if any pump parameters are beyond user-defined thresholds for more than two consecutive measurements (occurring every 5 s). Additionally, sound and light alarms would be triggered and notifications sent to system managers.
@@ 4.4 DAQ software
Each ARC controller communicates via an optical fiber to the DAQ computer through the custom-built CCDDrone [60] software, which sends configuration files, executes the readout sequence and receives the output images. Both are controlled centrally by the ccd-cdaq [61] Python library, designed to control CCD experiments. It loads another Python library, libabcd [62], which implements the MQTT protocol to communicate between processes, and logs, for debugging purposes, all messages that pass through the MQTT broker. A set of four applications (run control, DAQ status monitor, CCD manager, and spy) that run continuously in the background send and receive messages in an orderly fashion. Two CCD client applications, one for each CCD controller, communicate with each CCDDrone program using a custom Python library.
@@
New firmware and software have been written for the ACM and the ccd-cdaq is used to operate the ACM and control the CCDs. This new system runs in the LBC since the installation of the ACM, see Sec. 4.1.
@@ 4.5 Data quality monitor
The DQM framework provides tools to generate, store and visualize histograms and scalar elements (monitoring elements), using MongoDB as the database backend. The first step is low-level data processing of raw images to obtain the charge value for pixels, calculate median baselines, resolution, dark current per row, etc. The time evolution of these parameters is monitored, for example, the stability of dark current which can be used to identify new hot columns. Standardized algorithms perform statistical tests on the monitoring elements and automate the data certification processes. Additionally, the DQM framework enables visualization of monitoring results, retrieval of DQM quantities from the conditions database, and dataset selections for physics analyses.
@@ 5 Radioactive backgrounds
This section describes our effort to control and reduce the radioactive background in the LBC. Radio-assay analyses were performed for materials used in the detector – particularly those installed inside the cryostat. Results of our own or public radio-assay measurements have been utilized in dedicated Monte Carlo simulations of the whole detector and served us to guide the development and construction phases of the LBC. We present event rates measured with two setups in the LBC and compare them with simulations. The original configuration of the LBC is called Setup 1 and it includes two 6k×$\times$4k CCDs in the CCD box with OFHC copper lids. Upgraded detector is called Setup 2 and consists of two CCD modules, each with four 6k×$\times$1.5k CCDs, and EFC lids.
@@ 5.3 Simulation
From Setup 1 to Setup 2, we have reduced the total background by 5.5 dru according to simulations. The most significant reduction is from upgrading the lids in the CCD box to EFC. Another major decrease from exchanging 6k×$\times$4k CCDs with two CCD modules. The predicted decrease can be compared with measurements shown in the last row in Table 2 and discussed in Sec. 5.4.
@@
Table 2: The background rates in dru for two setups used in the LBC. Rates are the average calculated for or measured by the top and bottom CCDs. In simulations, cold copper includes all copper parts but the CCD box. External gammas and neutrons were not included in simulations. The background measurements are described in Sec. 5.4.

### v2

@@ 2 Prototype Objectives
First results on dark matter-electron scattering [44, 24], including a search for a daily modulation in the event rate [45, 31], were published, and studies of alternative low-mass dark matter candidates are forthcoming. In addition, three doctoral theses describe work on the LBC [46, 47, 48].
The LBC went through various configurations and two of them are highlighted in this paper, particularly when we describe the detector design in Sec. 3 and achieved improvement of the radioactive background in Sec. 5. Our adopted nomenclature for these two setups is as follows: The first one is Setup 1 and it includes two CCDs, each of the pixel format 6k×$\times$4k, placed in the CCD box fully made from oxygen-free high-conductivity (OFHC) copper. Setup 2 consists of two CCD modules, each with four 6k×$\times$1.5k-pixel format CCDs and one flex cable, in the CCD box with lids made from electro-formed copper (EFC). Let us stress that the two setups were using commercial CCD controllers which were exchanged for low-noise custom-made readout electronics developed for DAMIC-M and briefly mentioned in Section 4.1.
@@ 3 Detector Description
The LBC is designed to provide $O(\mathrm{1})$ dru background environment for the operation of CCDs with modular electronics. The innermost region of the detector consists of a cold copper housing, which simultaneously serves to cool the CCDs and provide an infrared shield from the surrounding room temperature materials. This cold copper housing is enclosed by 2 cm of ancient Roman lead from a sunken galley [49, 38], then an additional 4-10 cm of very low background ("très faible activité" or TFA) lead. Kinked channels provide a pathway for the two-layer polyimide flex cables, used for CCD operation and readout, to exit the inner region eliminating a line-of-sight to the inner copper housing. Cold OFHC [^2] In reality, it is from OFE-OK copper grade by Luvata in Finland. Because it is comparable to oxygen-free high thermal conductivity copper used in other experiments, we will use acronym OFHC. copper "fingers" travel through four narrow gaps in the interior (warm) lead shielding with $\sim$1 mm clearance to provide thermal contact and mechanical support for the inner copper housing.
Outside of the interior lead shielding are four axially-spaced intermediate boards, which provide the first stage amplification to the CCD video signal. The boards connect to flex cables through commercial 50-pin connectors by TE Connectivity on each side. The output side of each board connects to a 45 cm long, five-layer CCD flex cable that exits below the lead shielding to commercial 50-pin D-Sub vacuum feedthroughs by Accu-Glass below.
@@ 3.1 Charge-Coupled Devices
These 6k×$\times$4k CCD have two skipper amplifiers on one side and two standard amplifiers on the opposite side. Unlike conventional CCD amplifiers, skipper amplifiers can be configured to make multiple, non-destructive charge measurements (NDCMs) [25, 26, 27, 28]. Skipper readout essentially moves the charge contained in each pixel back-and-forth into the readout node allowing for many measurements of the same pixel so that they can be averaged. Since the measurements are uncorrelated, the readout noise is reduced to $\sigma_{1}/\sqrt{N_{\rm{skip}}}$, where $\sigma_{1}$ is the single-sample readout noise (the standard deviation of a single charge measurement) and $N_{\rm{skip}}$ is the number of NDCMs. By taking a large enough number of NDCMs, the readout noise can reach the sub-electron level and single charges [^4] For the sake of clarity, we use the term electrons to indicate charge carriers detected in the CCD. However, holes are held in the pixels of p-channel LBNL CCDs. are resolved. The single-electron resolution also provides a straightforward way of calibrating the energy response of the detectors [28].
A two-layer CCD flex cable carries the clock voltages to and video outputs from each CCD. The cable is glued to an extended non-functional area of the CCD. To improve adhesion, rectangular silicon shims were glued between the CCD die and the flex. The glue is low-background Epotek 301-2 previously selected for DAMIC at SNOLAB. The cable lines are wedge bonded with 25-$\mu$m thick Al wire to corresponding pads on the CCD. Special electrostatic discharge (ESD) safety precautions, described in Sec. 3.6, have been implemented during handling, packaging and testing CCDs to reach high packaging yield. During this time, care is taken to minimize the exposure of the CCDs to the air and hence, the accumulation of radon progeny on the CCD surface.
@@ 3.2 Boards and Cabling
A second-stage 45 cm long flexible cable, still thermally isolated from warm lead, goes from the intermediate board down to the cross part of the cryostat, where it is plugged into a two-to-one board. This board merges two cables coming from opposite sides of the same CCD and connects to the vacuum feedthrough. In early 2023, when two CCD modules replaced the 6k×$\times$4k CCDs, new intermediate boards were installed. Because each CCD module has only one cable to carry all signals, previously used two-to-one boards were taken out.
@@ 3.3 Cold Copper Parts and IR Shielding
The CCDs are housed in a copper box which cools the devices and acts as an IR shield. The dominant thermal path to the CCDs is through the CCD flex cables which are clamped to the CCD box via t-bars. In this configuration, the temperature remains stable when the on-chip CCD amplifiers are powered as detailed in Sec. 6.1.
@@
The connection to the cold tip of a cryocooler is made by a commercial flexible thermal strap by Thermal Management Technologies. Individual copper pieces were tightly connected with brass screws and a thin layer of thermally conductive grease (i.e. Apiezon N cryogenic high vacuum grease) was applied between them. For an electrical isolation, two 25 $\mu$m-thick Kapton tapes are placed between the cold tip and flexible strap.
@@
Because the copper lids are directly above and below CCDs, they are a major source of radioactive background. OFHC copper lids were used while taking science data with Setup 1 in 2022 [24, 31], but were taken out later the same year. The new installed lids are from less radioactive, electro-formed copper (EFC) grown and machined by Laboratorio Subterraneo de Canfranc (LSC) [56]. The EFC lids not only have much lower radioactive contamination than the commercial OFHC copper, but also their cosmogenic activation time has been reduced from many months to less than a week as described in Sec. 5.1.
@@ 3.5 External Passive Shielding and Support Structure
External gamma radiation and neutrons are attenuated by a layer of low background lead (faible activitè or FA) from LEMER and HDPE by Abaqueplast, both in France. The shielding around and above the cryostat cylinder, where the shielding of CCDs is the most important, consists of 15 and 20 cm of lead and HDPE, respectively, see Fig. 1 and 5. The bottom cross of the cryostat is surrounded by 5 cm of lead from all sides but below the cryostat and 10 cm of HDPE. The 41 cm diameter ring of lead shielding above the cryostat together with HDPE half-disks is supported by a 2 cm thick stainless steel plate.
The support structure standing on the cleanroom floor is static while a motorised opening system can move half of the shield on rails, allowing easy access to the cryostat. Neither the motor nor rails use lubrication and thus do not compromise the cleanroom cleanliness. A photo of the fully assembled and partly opened detector is in Fig. 5.
@@ 3.6 LSM Infrastructure
The furniture and equipment inside the cleanroom are ESD safe to comply with requirements on handling ESD-sensitive equipment, like CCDs and electronics. While working with CCDs, we require the humidity above 35% and we deploy a local humidifier using deionized water if needed.
The radon concentration has been continuously measured every 10 minutes with Durridge RAD7 in the cleanroom since January 2023. The median radon concentration is (18.4$\pm$8.5) Bq/m3 with 99% of the data below 42 Bq/m3. Other backgrounds were measured by different groups as part of the characterisation of the radiation environment in the laboratory, and they are: gamma flux of 3.8$\times$10-6 cm-2s-1 between 4 and 6 MeV [58], thermal neutron flux of $4.6\times 10^{-6}$ n/cm2/s at the LBC location [59], and fast neutron flux above 1 MeV of (1.06$\pm$0.60)$\times$10-6 n/cm2/s [60]. Details on these measurements can be found in given references.
@@ 4 Detector Readout and Instrumentation
The LBC is remotely controlled and monitored by custom data acquisition systems placed outside the external shield. The LBC is remotely controlled and monitored by custom data acquisition systems placed outside the external shield. We describe two different electronics systems installed in the LBC to run CCDs, of which the first one collected the data published in two DAMIC-M publications [24, 31]. The slow control, DAQ software and data quality monitor system are also detailed.
@@ 4.1 Readout Electronics
A custom data acquisition system based on a commercial Gen-III CCD controller from Astronomical Research Cameras, Inc. (ARC) was used for initial tests with the LBC and to collect data used in two papers [24, 31]. The controller provides the voltage biases and clocks required for operation of the CCD and reads out the video signal from two CCD amplifiers. External low-noise power supplies provide DC voltages for controllers, amplifiers and Vsub. Two CCD controllers were used to read two 6k×$\times$4k CCDs or two CCDs in two CCD modules.
@@
The new DAMIC-M readout electronics were successfully installed in the LBC in January 2024 and are used to control CCDs since then. These electronics were developed for DAMIC-M, have lower readout noise than the ARC CCD controller and can read all four CCDs in a module. The electronics chain consists of a custom-made Acquisition and Control Module (ACM) plugged to a VME crate (Wiener VME 6023x610), a 3 m long cable by Samtec and a custom-made front-end board. In addition, an adapter board connects the 3 m cable and D-Sub-50 connector in the vacuum feedthrough. More details on the new electronics will be given in a separate publication.
@@ 4.2 Instrumentation
The temperature of the cold copper is monitored by resistance temperature detectors (RTDs) PT-103 epoxied in cold fingers and cold mass table, see Fig. 3. The cooling rate and temperature are set by a Lakeshore 336 controller that monitors the RTDs and controls the power to a 100 W heater installed 2.5 cm next to the RTD closest to the cryocooler. To minimize temperature-induced stress on CCDs, we use the cooling rate of 0.1 K/min during temperature cycles. The system maintains the temperature within $\pm$0.2 K for many months, see Sec. 6.2.
@@
The detector and its components are protected by two-levels of backup power: (1) the DAMIC-M APC 2200 UPS and (2) the LSM system-wide UPS. In the case of a power outage, both UPS systems would be able to operate from battery power for at least one hour and supply instruments as shown in Fig. 6. This is adequate time for remote shifters to respond to alarms (see Sec 4.3) and safely shut down the experiment. In the unlikely scenario of an outage longer than one hour with no intervention from shift scientists, instruments would undergo a hard shut down. We have verified that the transition to UPS power is smooth and maintains the SC and DAQ systems up in the case of a power cut.
@@ 4.3 Slow Control System
Conditions that can pose a hazard for the operation of the CCDs are triggered on by alarm daemons. For example, a failure of the vacuum pump would cause moisture to condense on the cold CCDs and introduce shorts in the electrical components. To prevent such a scenario, the pump is powered by the local UPS which is monitored by the slow control system as is shown in Fig. 6. These alarms signal to on-shift users that the CCDs must be promptly shut down. In addition, a safety watchdog daemon is used to execute a shutdown if any pump parameters are beyond user-defined thresholds for more than two consecutive measurements (occurring every 5 s). Additionally, sound and light alarms would be triggered and notifications sent to system managers.
@@ 4.4 DAQ software
Each ARC controller communicates via an optical fiber to the DAQ computer through the custom-built CCDDrone [62] software, which sends configuration files, executes the readout sequence and receives the output images. Both are controlled centrally by the ccd-cdaq [63] Python library, designed to control CCD experiments. It loads another Python library, libabcd [64], which implements the MQTT protocol to communicate between processes and logs it. For debugging purposes, all messages that pass through the MQTT broker. A set of four applications (run control, DAQ status monitor, CCD manager, and spy) that run continuously in the background send and receive messages in an orderly fashion. Two CCD client applications, one for each CCD controller, communicate with each CCDDrone program using a custom Python library.
@@
The described DAQ software is used also for new ACM controllers installed in the LBC recently, see Sec. 4.1. To achieve this, we have written new firmware and modified local DAQ software.
@@ 4.5 Data quality monitor
The DQM framework provides tools to generate, store and visualize histograms and scalar elements (monitoring elements), using MongoDB as the database backend. The first step is low-level data processing of raw images to obtain the charge value for pixels, calculate median baselines, resolution, dark current per row, etc. The time evolution of these parameters is monitored. For example, the stability of the dark current can be used to identify new hot columns. Standardized algorithms perform statistical tests on the monitoring elements and automate the data certification processes. Additionally, the DQM framework enables visualization of monitoring results, retrieval of DQM quantities from the conditions database, and dataset selections for physics analyses.
@@ 5 Radioactive backgrounds
This section describes our effort to control and reduce the radioactive background in the LBC. Radio-assay analyses were performed for materials used in the detector – particularly those installed inside the cryostat. Results of our own or public radio-assay measurements have been utilized in dedicated Monte Carlo simulations of the whole detector and served us to guide the development and construction phases of the LBC. We present event rates measured with two setups in the LBC and compare them with simulations. The original configuration of the LBC, called Setup 1, includes two 6k×$\times$4k CCDs in the CCD box with OFHC copper lids. The upgraded detector, Setup 2, and consists of two CCD modules, each with four 6k×$\times$1.5k CCDs, and EFC lids.
@@ 5.3 Simulation
From Setup 1 to Setup 2, we have reduced the total background by 5.5 dru according to simulations. The most significant reduction is from upgrading the lids in the CCD box to EFC. Another major decrease comes from exchanging 6k×$\times$4k CCDs with two CCD modules. The predicted decrease can be compared with measurements shown in the last row in Table 2 and discussed in Sec. 5.4.
@@
Table 2: The background rates in dru (events/kg/kev/day) for two setups used in the LBC. Provided rates are the mean calculated for the top and bottom CCDs. In simulations, cold copper includes all copper parts but the CCD box. External gammas and neutrons were not included in simulations. The background measurements are described in Sec. 5.4.